# Efficient Backdoor Attacks for Deep Neural Networks in Real-world Scenarios

### v1

@@ 5.3 RQ3: Are proposed technologies stealthy for victims?
Fig. 8 showcases examples of poisoning images generated by different attacks on the ImageNet-50******As shown in Appendix 7.1, CIFAR-10 and CIFAR-100 have low resolution, which makes unclear visualizations. Therefore, we show the results on the ImageNet-50 dataset in this section. dataset. While our CLIP-UAP and CLIP-CFA may not achieve the highest stealthiness in terms of SSIM (as indicated in Table 2), the poisoning images generated by our methods appear more natural to human inspection compared to the baseline attacks. Additionally, incorporating CLIP-CFE has minimal impact on both PSNR and the natural appearance of the images, while achieving higher stealthiness in terms of SSIM.
Experiments on different poison rates for number-constrained backdoor attacks. We conduct ablation studies to assess the effectiveness of our proposed methods in reducing the number of poisoning samples (poisoning rates) for number-constrained backdoor attacks. The results depicted in Fig. 9 demonstrate the following: i) The attack success rate increases with higher poisoning rates for different attacks. ii) Our proposed CLIP-UAP and CLIP-CFA outperform the baseline techniques, BadNets [16] and Blended [5]. iii) The incorporation of our proposed CLIP-CFE further enhances the poisoning effectiveness across different triggers.
Experiments on different poison classes for class-constrained backdoor attacks. We conduct ablation studies to assess the effectiveness of our proposed methods in increasing the number of poisoning classes for class-constrained backdoor attacks. The results presented in Fig. 10 demonstrate the following: i) The attack success rate increases with higher poisoning classes for different attacks. ii) The attack success rate of clean-label single-class attack is lower than that of dirty-label single-class attacks. iii) Our proposed methods, CLIP-UAP and CLIP-CFA, outperform the baseline techniques, BadNets [16] and Blended [5]. iv) The incorporation of our proposed CLIP-CFE further enhances the poisoning effectiveness across different triggers.
Experiments on different domain rates for domain-constrained backdoor attacks. We conduct ablation studies to assess the effectiveness of our methods in increasing the domain rate for domain-constrained backdoor attacks. The results depicted in Fig. 11 demonstrate the following: i) The ASR increases with higher domain rates for different attacks. ii) Our proposed CLIP-UAP and CLIP-CFA outperform the baseline techniques, BadNets [16] and Blended [5]. iii) The incorporation of our proposed CLIP-CFE further enhances the poisoning effectiveness across different triggers.
Experiments on different large pre-trained models. We utilize the pre-trained CLIP model as the basis for our technologies. It’s worth noting that the community has proposed various CLIP variants. Therefore, an important practical consideration is whether our proposed technologies remain robust when applied different pre-trained CLIP models. To investigate this, we conduct ablation studies on different CLIP models for number-constrained backdoor attacks, as depicted in Fig. 12. The results demonstrate that our proposed technologies exhibit robustness across different CLIP models, with ViT-B/32 emerging as a competitive choice for all methods.
Figure 9: The attack success rate on the CIFAR-100 dataset with different poisoning rates. All results were computed as the mean of five different runs.
Figure 10: The attack success rate on the CIFAR-100 dataset with different accessible class of poisoning samples, where 1 (0) and 1 (1) in the abscissa represent the clean-label and dirty-label single-class attacks, respectively. All results were computed as the mean of five different runs.
Figure 11: The attack success rate on the CIFAR-100 dataset with different domain rates. All results were computed as the mean of five different runs.

### v2

@@ A.7.1 Datasets
A.7.2 Model Architecture.
We verify the performance on three popular model architectures of image classification: VGG-16 Simonyan & Zisserman (2014), ResNet-18 He et al. (2016), and MobileNet-V2 Sandler et al. (2018). All of them are widely used in various areas of artificial intelligence, such as flower classification Xia et al. (2017), pulmonary image classification Wang et al. (2019b), fault diagnosis Wen et al. (2020), and Covid-19 screening Farooq & Hafeez (2020).
A.7.3 Baseline and Comparison.
Our method contains two aspects: clean feature suppression and poisoning feature augmentation. Among them, poisoning feature augmentation can be accomplished through designing efficient and data-independent triggers, while clean feature suppression is orthogonal to previous trigger designing and can be integrated into any backdoor triggers. Therefore, we compare our two designed triggers CLIP-based universal adversarial perturbations (CLIP-UAP) and CLIP-based contrastive feature augmentation (CLIP-CFA) with two popular triggers: BadNets Gu et al. (2019)and Blended Chen et al. (2017). All of them are independent of the training data therefore can be implemented easily in the introduced data-constrained backdoor attacks. Although there are also contain other advanced clean-label backdoor attacks Liu et al. (2020); Barni et al. (2019); Saha et al. (2020); Souri et al. (2022) and state-of-the-art backdoor attacks Zhong et al. (2020), a substantial proportion of them Liu et al. (2020); Saha et al. (2020); Souri et al. (2022); Zhong et al. (2020) operate within the threat model that necessitates a proxy model pre-trained on the entire training set. This precondition becomes challenging to fulfill in the context of a many-to-one (M2O) data collection attack scenario. To verify the validity of the clean feature suppression, we integrate the proposed CLIP-based clean feature erasing (CLIP-CFE) onto currently designed triggers: two our designed triggers and two baseline triggers.
@@
A.7.4 Implementations.
In order to demonstrate the effectiveness of our proposed method, we conduct experiments on three datasets (CIFAR-10, CIFAR-100, and ImageNet-50). For the CIFAR-10 and CIFAR-100 datasets, we choose VGG-16, ResNet-18, and MobileNet-V2 as the victim models. All models use the SGD optimizer with a momentum of 0.9, weight decay of 5e-4, and a learning rate of 0.01 (0.1 for MobileNet-V2), which is multiplied by 0.1 at epoch 35 and 55. For the ImageNet-50 dataset, we use VGG-16 and MobileNet-V2 as the victim models. We use the SGD optimizer with a momentum of 0.9, weight decay of 5e-4, and a learning rate of 0.05 (0.01 for VGG-16), which is multiplied by 0.1 at epoch 35 and 55. The complete training epochs is 70.
In the number-constrained scenario, we conducted experiments with poisoning rates of 0.01 (P=500), 0.015 (P=750), and 0.007 (P=453) for the CIFAR10, CIFAR100, and ImageNet-50 threes datasets, respectively. In the class-constrained scenario, experiments with poisoning rates of 0.02 (P=1000), 0.01 (P=500), and 0.02 (P=1296) for three datasets. We choose two extreme scenario in the class-constrained backdoor attacks, denoted as clean-label single-class backdoor attack and dirty-label single-class backdoor attack. Specifically, the accessed class category is set to $Y^{\prime}=\{k\}$ and $Y^{\prime}=\{c\},c\neq k$ for clean-label single-class backdoor attack and dirty-label single-class backdoor attack, respectively. In the domain-constrained scenario, experiments with poisoning rates of 0.02 (P=1000, 1000, and 1296) for three datasets. The out-of-domain samples of all experiments are selected from other ImageNet-1K datasets that are not ImageNet-50. The attack-target class $k$ is set to category 0 for all experiments of above three data-constrained backdoor attacks.
A.7.5 Evaluation Metrics.
We evaluate the performance of our method in terms of Harmlessness: Benign Accuracy (BA), Effectiveness: Attack Success Rate (ASR), and Stealthiness: Peak Signal-to-noise Ratio (PSNR) Huynh-Thu & Ghanbari (2008) and Structural Similarity Index (SSIM) Wang et al. (2004).
Benign Accuracy (BA). BA is the clean accuracy of the testing set $\mathcal{D}_{t}=\{(x_{i},y_{i})|i=1,\cdots,M\}$ and is applied to evaluate the Harmlessness of the backdoor. When BA of the infected model is similar to the accuracy of the clean model, we believe that the current attack technique is harmless.
Attack Success Rate (ASR). ASR is applied to evaluate the effectiveness of the backdoor attack, which is the fraction of testing images with specific trigger that are predicted as the target class. Specifically, For $M^{\prime}$ images in the testing set that do not belong to the attack-target class ($k$), the ASR is formulated as:
$\text{ASR}=\frac{\sum_{i=1}^{M^{\prime}}\mathbb{I}(f(\mathcal{T}(x_{i},t);\Theta)=k)}{M^{\prime}},\quad(x_{i},y_{i})\in\mathcal{D}^{\prime}_{t},$ (15)
where $\mathcal{D}^{\prime}_{t}$ is a subset of testing set $\mathcal{D}_{t}$ ($\mathcal{D}^{\prime}_{t}\subset\mathcal{D}_{t}$), containing the images whose label is not the attack-target class $k$.
Peak Signal-to-noise Ratio (PSNR) Huynh-Thu & Ghanbari (2008). PSNR is applied to measure the similarity between clean images and the corresponding poisoning images. Give a image $x_{i}\in\mathcal{D}_{t}$ and the corresponding poisoning image $x^{\prime}_{i}=\mathcal{T}(x_{i},t)$, the PSNR is formulated as:
$\operatorname{PSNR}=\frac{1}{M}\sum_{i=1}^{M}\operatorname{PSNR_{i}}(x_{i},x^{\prime}_{i}),$ (16)
$\operatorname{where}\quad\operatorname{PSNR_{i}}(x_{i},x^{\prime}_{i})=10\log_{10}\left(255^{2}/\operatorname{MSE}(x_{i},x^{\prime}_{i})\right),$
$\operatorname{MSE}(f,g)=\frac{1}{HW}\sum_{i=1}^{H}\sum_{j=1}^{W}\left(f_{ij}-g_{ij}\right)^{2},$
$H$ and $W$ are height and width of the image, respectively. Larger PSNR means larger similarity between clean images and the corresponding poisoning images, therefore larges stealthiness of backdoor attacks.
Structural Similarity Index (SSIM) Wang et al. (2004). Similar to PSNR, SSIM is another metrics to represent the stealthiness of backdoor attacks, which is formulated as:
$\operatorname{SSIM}=\frac{1}{M}\sum_{i=1}^{M}\operatorname{SSIM_{i}}(x_{i},x^{\prime}_{i}),$ (17)
$\operatorname{where}\quad\operatorname{SSIM_{i}}(x_{i},x^{\prime}_{i})=l(x_{i},x^{\prime}_{i})\cdot c(x_{i},x^{\prime}_{i})\cdot s(x_{i},x^{\prime}_{i}),$
$\left\{\begin{array}[]{l}l(f,g)=\frac{2\mu_{f}\mu_{g}+C_{1}}{\mu_{f}^{2}+\mu_{g}^{2}+C_{1}}\\ c(f,g)=\frac{2\sigma_{f}\sigma_{g}+C_{2}}{\sigma_{f}^{2}+\sigma_{g}^{2}+C_{2}}\\ s(f,g)=\frac{\sigma_{fg}+C_{3}}{\sigma_{f}\sigma_{g}+C_{3}}\end{array},\right.$
where $\mu$ and $\sigma$ are mean and variance of image, respectively. Similarly, Larger SSIM means larger similarity between clean images and the corresponding poisoning images, therefore larges stealthiness of backdoor attacks.
Figure 9: Defense results of pruning Liu et al. (2018) on the number-constrained backdoor attack of the CIFAR-100 dataset and VGG-16 model, where the number of clean samples owned by the defender is 50 and the poisoned rate is $2\%$.
Figure 10: Defense results of Neural cleanse Wang et al. (2019a) on the number-constrained backdoor attack of the CIFAR-100 dataset and ResNet-18 model, where the number of clean samples owned by the defender is 1500, and the poisoned rate is $2\%$.
Figure 11: Defense results of FST Min et al. (2023) on the number-constrained backdoor attack of the CIFAR-100 dataset and ResNet-18 model, where the number of clean samples owned by the defender is from 1000 to 2500 and the poisoned rate is $2\%$.